Express 5 cm in meter and kilometer
 Express 5 cm in meter and kilometer

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

We know that,
= 1 meter = 100 cm
Then,
= 1 cm = (1/100) m
= 5 cm = (5/100)
= 0.05 m
Now,
= 1 km = 1000 m
Then,
= 1 m = (1/1000) km
= 0.05 m = (0.05/1000)
= 0. 00005 km

Thanks